OpenStudy (buttermequeasy):

Take the points you have selected and write the equation of the line using the point-slope formula. Change the equation to slope-intercept form. Show the work. (2,5) and (3,4)

OpenStudy (buttermequeasy):

Help please I don't understand!

OpenStudy (anonymous):

use the slope formula m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

then plug in the value for m in the equation y-mx+b

OpenStudy (anonymous):

solve for b

OpenStudy (buttermequeasy):

That's what I don't understand.

OpenStudy (anonymous):

btw, I meant y=mx+b

OpenStudy (anonymous):

m=(4-5)/(3-2) m=-1

OpenStudy (anonymous):

sub point (2,5) and m=-1 into y=mx+b and solve for b

OpenStudy (moongazer):

How do you solve for b?

OpenStudy (anonymous):

5=(-1)(2)+b b=7 thus the equation of the line is y=-x+7